# Empirical Analysis Of Heuristic and Approximation Algorithms for the The Mutual-Visibility Problem

### v2

@@ 1. Introduction
The mutual-visibility (MV) problem was introduced by Di Stefano et al. in (Di Stefano, 2022) and was subsequently generalized and extended in (Cicerone et al., 2023). Given a subset $X\subseteq V(G)$, we say that vertices $u,v\in X$ are X-visible if there exists a shortest path $P$ such that $V(P)\cap X\subseteq\{u,v\}$. Such a set $X$ is called a mutual-visibility set if every pair $u,v\in X$ are X-visible.
@@ 2.1. Direct Random Algorithm
[…]
@@ 2.2. Genetic Algorithm (GA)
The general structure of the adapted GA is outlined in Algorithm 2.
Input: Graph $G$; maximum iterations $T$; population size $N$; crossover probability $p_{c}$; mutation probability $p_{m}$; penalty parameter $M$
@@
1 Initialize a population $\mathcal{P}$ of $N$ individuals
@@ 8       Evaluate fitness of all individuals in the new population
9       Select the best $N$ individuals to form the next generation $\mathcal{P}$
@@ 8       Remove all edges incident to $y^{*}$ from $E(H^{\prime})$
Mycielskian of Paths ($M(P_{n})$) Mycielskian operation on path graphs. $n$ (base) $n+\lfloor\frac{n+1}{4}\rfloor$ for $n\geq 5$ (and $6$ for $n=4$)
Mycielskian of Star Graphs ($M(K_{1,k})$) Mycielskian on stars (base has universal vertex). $k+1$ (base nodes) $12k+$ 1 for $k\geq 1$ (from $12n-$ 1 for base $n$)
@@ 4.1. Overall Performance Overview
We first provide a high-level summary of the algorithm’s average performance for Category 1 of the dataset. Notice that in larger graph instances, the average performance of GA can occasionally suffer due to the algorithm getting stuck in local maxima. This results in a few individual cases exhibiting very poor approximation ratios, while other individual runs on similar instances might still perform well. These findings are visible in Tables 2, where $\overline{\alpha}$ represents average ratio for the approach, and 3, where $\overline{T}$ represents average runtime for the approach.
@@ 4.2. Approximation Quality Analysis
[…]

### v3

# Empirical Analysis Of Heuristic And Approximation Algorithms For The Mutual-Visibility Problem

@@ 1. Introduction
The mutual-visibility (MV) problem was introduced by Di Stefano et al. in (Di Stefano, 2022) and was subsequently generalized and extended in (Cicerone et al., 2023). Given a subset $X\subseteq V(G)$, we say that vertices $u,v\in X$ are X-visible if there exists a shortest path $P$ such that $V(P)\cap X\subseteq\{u,v\}$. Such a set $X$ is called a mutual-visibility set if every pair $u,v\in X$ are X-visible. A $\mu$-set is a maximum mutual-visibility set (i.e., no larger mutual-visibility set exists). When we only require that no proper superset is an mutual-visibility set, we speak of a maximal MV-set. Let $\mu(G)$ denote the cardinality of a $\mu$-set of graph G.
@@ 2.1. Direct Random Algorithm
We construct a simple random algorithm for finding the largest mutual-visibility set within a graph $G$. The algorithm works by random sampling; by first choosing a subset size $k$ uniformly from $1..n$, then sampling a $k$-subset of $V(G)$, and checking for the mutual-visibility property of the set. As a default the number of trials $T$ was set to $10,000$. The algorithm is described in Algorithm 1. Note that the procedure $\text{MV}(G,S_{candidate})$ is defined in (Di Stefano, 2022, p. 7) as Procedure MV.
@@ 7     $S^{*}\leftarrow S$
The general structure of the adapted GA is outlined in Algorithm 2. The parameters $p_{c}$ and $p_{m}$ were both chosen simply as $1.0$.
Input: Graph $G$; maximum iterations $T$; population size NN; crossover probability $p_{c}$; mutation probability $p_{m}$; penalty parameter $M$
@@
1 Initialize a population $\mathcal{P}$ of NN individuals
@@ 8  Evaluate fitness of all individuals in the new population
9  Select the best NN individuals to form the next generation $\mathcal{P}$
@@ 8  Remove all edges incident to $y^{*}$ from $E(H^{\prime})$
Mycielskian of Paths ($M(P_{n})$) Mycielskian operation on path graphs. $n$ (base) $n+\lfloor\frac{n+1}{4}\rfloor$ for $n\geq 5$ (and 66 for $n=4$)
Mycielskian of Star Graphs ($M(K_{1,k})$) Mycielskian on stars (base has universal vertex). $k+1$ (base nodes) $2k+1$ for $k\geq 1$ (from 2n12n-1 for base $n$)
@@ 4.1. Overall Performance Overview
We first provide a high-level summary of the algorithm’s average performance for Category 1 of the dataset. Notice that in larger graph instances, the average performance of GA can occasionally suffer due to the algorithm getting stuck in local maxima. This results in a few individual cases exhibiting very poor approximation ratios, while other individual runs on similar instances might still perform well. These findings are visible in Tables 2, where $\overline{\alpha}$ represents average ratio for the approach, and 3, where $\overline{T}$ represents average runtime for the approach. Notice that the genetic algorithm has $\overline{\alpha}$ over 1 in some cases, this is due to its indeterministic nature, where it can choose a bigger $S$ than $\mu(G)$, and it does not optimize it in the specified maximum iterations.
@@ 4.2. Approximation Quality Analysis
Our empirical analysis reveals that for graphs with a small number of vertices (e.g., $n=10$), the obtained mutual-visibility set sizes consistently approximate or closely align with the established theoretical lower and upper bounds. However, for larger graph instances (eg. $n=100$), this relationship becomes less evident. The random‐sampling algorithm provides a natural baseline against which to compare more complex methods for the mutual-visibility problem. Empirical results show that, with T=10  [^0] T=10\,000 trials, it often yields near-optimal estimates of the mutual-visibility number $\mu(G)$ on small (e.g. $n=10$) and even on certain larger graphs of certain classes. However, as the order of the graph increases, the quality of the approximation deteriorates unless the number of random trials is increased proportionally. Another observation is that the lower bound obtained from the hypergraph approximation method is on average lower than the general lower bound provided by the largest degree and hence not very useful in providing tighter bounds. For the genetic algorithm approach the relation between the achieved $\mu$-set size vs. general lower bound ($\min\{n-d+1,n-c+3\}$) can be observed from the graph in Figure 2. Note that the blue coloured data points represent analysis on general Petersen graphs, while orange data points represent Erdős-Rényi graphs.